OpenStudy (anonymous):

The radius of a circle is multiplied by 10. Its area is _____ times bigger than the original circle.

OpenStudy (anonymous):

would it 100

Parth (parthkohli):

The ratios of the areas is simply the ratio of the sides/radius.

Parth (parthkohli):

Yep, 100.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thank you

Parth (parthkohli):

\( \color{Black}{\Rightarrow r = 1 \Longrightarrow Area = 3.14 }\) \( \color{Black}{\Rightarrow r = 10 \Longrightarrow Area = 3.14 \times 100 = 314}\)

Parth (parthkohli):

We can try these out with radii/sides too.

Parth (parthkohli):

If a ratio is given as 1:5, then just assume that r = 1 and r = 5, and find the areas respectively.

OpenStudy (anonymous):

The ratios for sides and areas differ. The ratio when something happens to side measures or lengths, the effect is squared in the area.\[Sides: \frac{a}{b} \rightarrow \frac{1}{10}\]\[Area: \frac{a^2}{b^2} \rightarrow \frac{1}{100}\]

OpenStudy (anonymous):

ok thank you both for the explanation

OpenStudy (anonymous):

np :)